OpenStudy (anonymous):

you are rolling a ball off a table. if the tabletop is 1.22 m above the floor and the golf ball hits the floor 1.52 m horizontally from the table, what was the initial velocity of the golf ball?

OpenStudy (shane_b):

First, calculate the time it takes the ball to fall to the floor. Since it has no initial vertical velocity it will simply be:\[h=\frac{1}{2}gt^2\]\[1.22m=\frac{1}{2}(9.8m/s^2)t^2\]\[t=0.499s\] Now that you know must time it was falling you can figure out its initial horizontal velocity:\[V_x=\frac{displacement}{time}=\frac{1.52m}{0.499s}=3.046 m/s\]

OpenStudy (shane_b):

Of course, that velocity is only in the x direction...it had 0 initially in the y direction.